GreenHab Report – April 28th

GreenHab Report

Name: Shawna Pandya

Date: 28Apr2018

Environmental control: Ambient with door open

Shade cloth: on

Fan: Used today

Average temperatures: 39.0℃

Low: 15.5℃

High: 52.2℃

Hours of supplemental light: None

Changes to crops: Less yellow with additional watering

Daily water usage for crops: 10 Gallons

Time(s) of watering for crops: Twice: Once at 0900 and again at 1700

Morning research observations: None

Changes to research plants: No.

Aquaponics: None

Narrative: Picked a few tomatoes. Assembled red wagon – the latest ‘Mars Rover’!

Support/supplies needed: None

Astronomy Report – April 27th

Astronomy Report
Name: Eric Shear
Crew: 193
Date: April 27, 2018

Sky Conditions: Clear

Wind Conditions: None

Observation Start Time: 1016

Observation End Time: 1100



One solar flare was discovered today, aimed right at us. It is a whitish spot next to a filament, as shown in the photo attached.

No sunspots or prominences could be spotted with the Skyris camera, despite using the tuners.

Photos were taken with the Skyris camera and the best image processed in Autostakkert, Registax, Photoshop, and PowerPoint to colorize and bring out contrast. The result is attached.

Objects Viewed: Sun

Problems Encountered: None

Science Report – April 27th

Science Report for 27 April 2018
Crew 193 – PHEnOM Gold Crew
Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: No locations were scouted during the EVA today. Emergency triage procedures were practiced so that a crew member might be safely transported to an evacuation location in the event of an incident.

2. Sunspot and Solar Flare Monitoring: Our CGA continued his task of monitoring the sun for sunspot activity. He found a solar flare in the direction of our planet and immediately alerted mission control to the danger.

3. Shortwave Texting on EVA’s: Hand signal communication was used during the emergency EVA practice today. Further practice is needed to achieve a smooth process.

4. GPS Route Measurements: No measurements today, as the EVA was at the hab location.

5. Waterless dish cleaning: Samples were taken from a dirty dish and a dirty dish that had been put under the UV light. There are minimal signs of bacterial growth from the first set of samples. It may take a couple more days to see signs of growth. Both sets of samples are now in the incubator in the hab.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today. Possible that the continuation of this study from Crew 192 will be scrapped.

7. Spacesuit Helmet Fogging: Neither crew experienced fogging of the masks today. Both crews used a defogging spray and non-scratch cloth on the masks prior to their EVAs today.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Analysis will be done as the mission continues.

Journalist Report – April 27th

Journalist Report – Earth-Date 27APR2018 / Mars Sol 5
Author: Doug Campbell

It’s been a tale of two Sols over the last 48 hours for MDRS Crew 193. Sol 4 was packed with two extensive exploration EVAs, a delicious meal and a hard fought battle at the card table.

On Sol 4, Team Canada (EVA Team 1) left the vast expanses of the Martian desert to explore the moon. The soft, powdery, grey dirt enabled the team to travel vast distances and climb tall mountains with the ease of Alan Shepard hitting a golf ball in 1/6 gravity. They walked over 9 kilometres (5.5 miles for our imperial friends) and documented the unique landscape with their strong photographic skills. EVA Team 2 explored deep into Candor Chasma and marveled at the beauty of the valley. Unfortunately, due to a framing error, the breathtaking pictures that were anticipated did not turn out.

In the evening, the crew astronomer/geologist introduced his colleagues to a card game called “Shanghai”. The rest of the crew is figuring out the rules as they play, but they are beginning to wonder if the astronomer is trying to hustle them for extra dessert rations!

Sol 5 was supposed to be a relaxing day for crew bonding and skill building. However, things to not always go as planned! The day began with training on emergency triage procedures. This would enable the crew to care for their colleagues should an accident occur on an EVA. The newly learned techniques were then practiced during an EVA outside of the hab.

Two teams practiced triaging patients and moving unconscious astronauts into safer positions when they are trapped in an EVA vehicle. Unexpectedly, these skills were put into immediate real usage as one crew member reported they were feeling dizzy during the EVA. The other crew members responded immediately and brought the crew member safely inside the hab for treatment. We are happy to report the crew member is doing well and that the training their teammates received this morning enabled them to act quickly and decisively to avoid further issues.

The afternoon was spent recuperating, working on projects and furthering our research. It was not exactly the day the crew had planned; however, it showed the strength and adaptability of the crew to deal with issues as they arise.

A quiet evening of eating good food and watching a movie is anticipated. Perhaps the cards will also be broken out again…

Greenhab Report – April 27th

Name: Anima Patil-Sabale
Date: 27 Apr 2018

Environmental control: Ambient with door open but only for few hours in the afternoon
Shade cloth: on
Fan: not used today

Average temperatures: 34.8 ℃
Low: 15.5℃
High: 50.2℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 11 Gallons

Time(s) of watering for crops: Twice: 10:15am and 5:15pm

Morning research observations: None

Changes to research plants: No.

Aquaponics: None

Narrative: None

Support/supplies needed: None

Operations Report – April 27th

SOL: 4

Name of person filing report: David Attig

Non-nominal systems: A couple sim suit chargers, and soldering iron.

Notes on non-nominal systems: A sim-suit charging cable broke yesterday, due to having cold solder joints. I attempted to repair, but the soldering Iron broke after heating briefly.

Generator (hours run): 9:30 pm to 9:00 am.

Solar— SOC 22 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 550 gallons.

Water (static) – 250 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134192.9

Toilet tank emptied: no
Date last emptied: April 26

ATV’s (Honda, 300, 350.1, 350.2, 350.3): Yes, 300

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0.01 Gallons

# Hours the ATVs were Used today: 0.1

Notes on ATVs: 350.2 is out of service, due to a previous issue.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: Yes
Hours: 34.2
Beginning charge: 100 %
Ending charge: 96 %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Eric made progress troubleshooting the beartooth devices by communicating with the manufacturer. The issue has not been resolved yet. Last night we waited too long to turn on the generator, SoC had dropped to 22%. This morning, it was only 26% after the generator running all night. It appears that the generator is not charging the batteries at all. SoC returned to 100% after a few hours of full sunlight. Mission Control (Shannon) is investigating further.

Summary of internet: We’re out, same as normal.
Summary of suits and radios: One piece suits nominal, still work to do on two piece suits. I changed some settings on radios we use during EVA (whisper function) to make the microphone more sensitive. Recieved new radios, opened one set. They are ready to use on EVAs. Demoted some older EVA radios to science dome, retired one radio from science dome and one from the green hab (both radios have issues).

Summary of Hab operations: Started sorting tools in toolbox, and radios/chargers. Went out and checked water tank, confirmed we are half way through it.

Summary of GreenHab operations: Heavily watered (8+ gallons) in morning at request of mission control.

Summary of ScienceDome operations: Dishwasher experiment proceeding, more samples taken today.

Summary of RAM operations: New radios retrieved from RAM airlock.

Summary of health and safety issues: During a local EVA to practice medical emergencies, we had one crewmember suffer from heat exhaustion. EVA was cut short, crewmember is feeling better.

Questions, concerns and requests to Mission Support: We would like some small cardboard boxes for organizing the HAB. They would also be good for moving material and supplies during EVAs (such as to/from RAM).
We would like some hang-tags or white cloth (gaffing?) tape to mark things as we organize the HAB.
Also, we would like some small paintbrushes (for detail work on sign), wet-wipes, and paper towels.

EVA Report – April 27th

EVA 10

EVA Gold Crew Members – Gold000 (CO Patil-Sabale) / Gold 003 (CE Attig) /
Gold004 (HSO Pandya)

EVA Objectives – Summary

  1. Medical Skills, Triage, Decision-Making and Evacuation on EVA – Completed
  2. EVA Comfort, Well-Being, Safety and Optimization Study / Raw data here
  3. Visor Fogging Study – Data collected / Raw data here


  • Was very hot out, kept EVA short (70-80min), stayed at base, checked in with crew members frequently

ATVs/Rovers Used

  • Opportunity – 100% start; moved 15 ft for medical skill simulation; 96% end after testing after EVA 11
  • ATV 300 – Moved 15ft for medical skills simulation

Safety Measures & Preparation

  • Sunscreen/Sun protection – Yes/Medical Exemptions for UV eye-protection provided
  • Adequate hydration – No
  • Adequate protection – Flightsuit/footwear/visor/glove precautions – Yes

Sol 5 Crew 193 EVA 10&11 Summary
Earth-Date 27APR2018

EVA 11

EVA Gold Crew Members – Gold001 (XO Campbell) / Gold 004 (HSO Pandya) /
Gold005 (CGA Shear)

EVA Objectives – Summary

  1. Medical Skills, Triage, Decision-Making and Evacuation on EVA – Completed
  2. EVA Comfort, Well-Being, Safety and Optimization Study / Raw data here
  3. Visor Fogging Study – Data collected / Raw data here


  • Preemptively cut short due to heat exhaustion in one crew member, after about 50min

ATVs/Rovers Used

  • Opportunity – 100% start; moved 15 ft for medical skill simulation; 96% end after testing after EVA 11
  • ATV 300 – Moved 15ft for medical skills simulation

Safety Measures & Preparation

  • Sunscreen/Sun protection – Yes/Medical Exemptions for UV eye-protection provided
  • Adequate hydration – No
  • Adequate protection – Flightsuit/footwear/visor/glove precautions – Yes

Sol Summary – April 27th

Crew 193 Sol Summary Report – 27APR2018

Sol 5

Summary Title: All Aboard the Sooooollllll-Train

Author: Dr. Shawna Pandya, HSO

Mission Status: Nominal like you wouldn’t believe. Bigly Nominal. My friends tell me they’ve never seen a mission more nominal.

Sol Activity Summary: Today we planned for a shorter set of EVAs in anticipation of the hot day and gusting winds, concentrating on medical skills and simulation (see below). For lunch, we feasted on a delicious Martian lunch of Jamabalaya rice, broccoli and cheese soup, and fish. The afternoon was then spent on testing new radios and equipment, ongoing science experiments, finishing the MDRS sign, and usual Hab operations. Notably, our crew astronomer did a stellar job of picking up a solar flare, leading to the declaration of a state of emergency from the ensuing radiation storm, and reinforcing the good work we are doing scouting and mapping potential emergency shelters while on EVA. The remainder of the day was spend planning out our remaining week on Mars.

Look Ahead Plan: Tomorrow will be a mandatory rest day with no EVAs planned in anticipation of the 30C high. We will plan out the rest of our time here at MDRS, including a school visit on Sol 11, May the 4th on Sol 12, and end-of-rotation operations on Sol 13.

Anomalies in work:

1. Paintbrushes were not small enough for lettering on MDRS sign.
2. Soldering iron is broken.

Weather: Clear skies in AM, extremely hot by mid-day.

Crew Physical Status: Affected crew member from EVA recovering nicely.

EVA: The crew spent the morning going over principles of wilderness medical triage, evacuation and critical decision-making in case of accident while on EVA. We then moved these skills outside in our spacesuits, simulating various medical emergencies, including extraction from a rolled ATV while maintaining C-spine precautions (easier said than done in a spacesuit!). Our second EVA was cut short due to heat exhaustion in one crew member.

Reports filed:
EVA Report
EVA Request
Astronomy Report

Science Report
Journalist’s Report

Reports to be filed:
Photo & Photo of the Day Report

Operations Report

Support Requested:
1. Smaller paint brushes for MDRS signs.

Commander Report – 12 April 2018

Commander, PHEnOM Gold Crew, MDRS #193

Mission Sol 4 (Earth Day April 26th, 2018)

Its a beautiful feeling to be on Mars. We have made a home away from home here and are starting to make a special bond with this red planet.

We have adapted to this new routine, with our Executive Officer and Health Science Officer setting of on their morning EVA, while Crew Engineer, Crew Geologist and Astronomer and me working around the hab in the greenhouse, musk observatory and on hab maintenance. The plants in the greenhouse are doing great, we harvested fresh greens today to make salad, soup and to add flavor to our pizzas. While our CE has been doing a great job of maintaining the hab and the vehicles, our CGA has been impressive with his work in the observatory.

The afternoon EVA turned into another adventure when we couldn’t find a road mapped by rovers and previous crews, as it was stormed out and there was no trace left of it. We resorted to the map and GPS coordinates, used our navigational skills and tracked it down. Our efforts were paid off when we reached the canyons at Candor Chasma and were treated to beautiful views. We scouted for emergency shelters, but all that terrain is really flat to have any possibility of emergency shelters, unless we hike down into the canyons. We have left signs to mark part of the road that was washed off, so that future crews are able to find it.

We have had a really busy week so far. Our research projects are coming along great and am happy with the progress we are making.
Planning on winding down the next couple of days as the weather is predicted to be extreme. Will spend time on some team trainings, hab projects and planning for the next week. The team is in great health and spirits and looking forward to the next few days of adventures, research and exploration. I am impressed with the camaraderie and team spirit each crew member has exhibited. Proud to be a part of this crew.

Copyright © The Mars Society. All rights reserved. | Main Site